OpenStudy (anonymous):

If a light has a wavelength of 550nm, what is the energy of one photon if this light?

OpenStudy (ookawaiioo):

E = h* v h= plank constant v = (3.00*10^8)/ (5.5x10^-7) plug into the equation....... and thank me later